OpenStudy (anonymous):

A population of bacteria is growing according to the function below, where t is in hours. How many hours will it take for the population to grow to 1,000 bacteria? Round your answer to the nearest integer. Do not include units in your answer. B(t) = 50 e2.5t

OpenStudy (saifoo.khan):

wht sign is there before "e2.5t"

OpenStudy (anonymous):

\[B(t)=50\times e^{2.5t}\]

OpenStudy (anonymous):

e^2.5t = 20 t= ln(20) / 2.5

OpenStudy (anonymous):

set \[B(t)=1000\] and solve for t via \[1000=50e^{2.5t}\] \[20=e^{2.5t}\] \[\ln(20)=2.5t\] \[t=\frac{\ln(20)}{2.5}\]

OpenStudy (anonymous):

damn elecengineer beats me to the punch again!

OpenStudy (anonymous):

hey satellite, can you help me